# Multi-mode Pulsations in AGB Stars: Insights from 3D RHD CO5BOLD Simulations

### v1

@@ 1 Introduction
Mode switching, where a star changes its dominant pulsation mode (defined as the mode with the highest amplitude), is another phenomenon closely tied to stellar evolution. This transition is expected as stars evolve across different P-L sequences, with observational evidence showing that red giants can switch between fundamental and overtone modes (Kiss et al. 2000; Hinkle et al. 2002). The combined amplitude of decaying higher-overtone modes and emerging lower-overtone modes may be a key factor in triggering more complex pulsation behaviours, such as the long secondary period (Pawlak 2021). The transition from small-amplitude red giants to semiregular variables (SRVs), associated with the onset of self-excited pulsations, is linked to changes in the stellar envelope (Trabucchi & Pastorelli 2025). These changes in the pulsation modes are further influenced by factors such as metallicity, hydrogen content, and the treatment of turbulent viscosity in the models. The sensitivity of the transitions of the pulsation modes to the stellar envelope composition provides critical constraints for theoretical models, particularly in understanding the evolutionary pathways of long-period variables (LPVs) (Trabucchi & Pastorelli 2025). Mode switching offers key insights into the internal structure of stars and the mechanisms governing the stability of pulsation modes, with broader implications for stellar feedback and the calibration of P-L relations.
@@ 2.2 Advances over previous generation of models
A significant improvement in this new grid is the adjustment and consistent application of the gravitational smoothing parameter, $r_{0}$ (see Ahmad et al. 2023), across the new model grid. This was set to $7\,\%$ of the stellar radius ($\mathrm{R}_{\star}$), in comparison to between $20-25\,\%$ of $\mathrm{R}_{\star}$ in our previous models. The stellar radius, $\mathrm{R}_{\star}$, is defined as the point of minimum entropy, based on prior investigations into the optimal definition of this parameter in our models (Ahmad et al. 2023). The $r_{0}$ parameter prescribes the smoothing of the stellar gravitational potential to account for the unresolved compact core, where most of the stellar mass resides, given that we have set the core mass of $0.6\,\mathrm{M}_{\odot}$. Within $2\,r_{0}$, energy is injected at a constant rate according to the desired luminosity. The size of the computational domain and the number of grid points were chosen to keep the number of cells per surface pressure height similar and to account for the different extensions of the atmospheres of models with different surface gravities.
@@ 2.3 Pulsation period extraction process
The extraction of pulsation periods begins with the application of the Fast Fourier Transform (FFT) to temporal quantities selected according to the specific characteristics of the pulsation modes under investigation. Unless explicitly stated otherwise, all references to power spectra in this article pertain to those derived from the FFT. To analyse purely radial motions, the radial mass flux and mass density are averaged over spherical shells. These spherically averaged quantities are denoted as ¡.¿Ω, where $\Omega$ represents the solid angle of a sphere. The ratio of the averaged radial mass flux, ¡$\rho v_{\textrm{radial}}$¿(${}_{\Omega}(r,t)$ to the averaged mass density, ¡$\rho$¿(${}_{\Omega}(r,t)$, yields the radial velocity as a function of radial distance and time, ¡$v_{\textrm{radial}}$¿$(r,t)$. This approach is consistent with the methods used in Freytag et al. (2017) and Ahmad et al. (2023). To account for non-radial pulsation modes, this method is extended using spherical harmonics. By incorporating weights associated with spherical harmonics into the spherical averaging, where each degree is denoted with $\ell$, there are $12\ell+$ 1 spherical harmonics ($-\ell\leq m\leq+\ell$); we obtain the velocity as a function of $\ell$ and $m$. For each degree, we sum the power spectra across each possible $m$ to obtain a single power spectrum representative of each $\ell$.
@@
Applying the FFT to the spherically averaged velocity yields FFT spectra as a function of radial distance from the stellar centre, corresponding to layers within the stellar interior. To account for spatial variability across these radial layers, an average power spectrum is computed for the region spanning $80-100\,\%$ of $\mathrm{R}_{\star}$. This region consistently exhibits significant contributions from pulsations and their effects. To enhance the clarity of signal detection, a combination of sequential smoothing methods is applied. First, a Savitzky-Golay filter is used, followed by convolution with a Gaussian kernel. This combined approach preserves the peaks in the power spectrum, while effectively reducing high-frequency noise. Overall, this ensured reliable identification of pulsation features in Fourier space, and mitigates non-linear effects such as interference by convection.
Significant frequencies are identified using an iterative peak detection algorithm. At each iteration, a region of considerable signal is isolated from the entire power spectrum. Within this region, the weighted mean frequency ($f_{\textrm{mean}}$), root mean square (RMS) values ($f_{\textrm{RMS}}$), and the total power contained within the region are normalised by the total power of the spectrum to calculate the power fraction, representing the significance of the signal. The identified region is then removed from the power spectrum, and the process is repeated. This procedure is analogous to the pre-whitening techniques in asteroseismology, where successive peaks are identified and removed to isolate additional modes. Up to 20 regions are identified in this manner, then, any overlapping regions are combined into single regions. For these combined regions, the mean, RMS, and power fraction are recalculated. Within these regions, the pulsation period and its spread (could be accounted for as its uncertainty) is then calculated with $P_{\textrm{puls}}=1/f_{\textrm{mean}}$, and its spread $\sigma_{\textrm{puls}}=P_{\textrm{puls}}(f_{\textrm{RMS}}/f_{\textrm{mean}})$.
@@
The method outlined above successfully identifies various pulsation modes across different radial orders. The overtone order is determined systematically: the fundamental mode corresponds to the lowest frequency, the first overtone to the second-lowest, and so on. Non-radial modes are distinguished by the velocity components used in the power spectrum analysis. The derived pulsation frequencies and their spreads are used to calculate the fraction of power within each mode’s frequency band. Consequently, we now have periods, frequency spreads, and power fractions for each radial order and degree, allowing for direct comparisons between each of the derived modes. To evaluate luminosity variations, we calculate the spherically averaged luminosity at the outer boundary of the computational domain. While spherical averaging may retain some imprints of non-radial modes, these are less distinct and not reliably distinguishable. However, radial orders remain well separated and can be confidently analysed.
@@ 2.4 Radial node extraction process
Figure 1: The phase-normalised amplitude derived from the Fourier transform, in the frequency bands of the first and second overtones, in the top and bottom row, respectively.
We further utilise the velocity data within the star to identify radial nodes when an overtone is detected using the process described above. First, we take the quantity $\sqrt{\rho}v_{\mathrm{radial}}$ as an effective measure of density variations within the star. The complex Fourier transform of this quantity is calculated, now as a function of radial distance and frequency, and subsequently used in the following steps.
@@
The process begins with phase normalisation, aligning the phase across the radial layers to the phase as computed at the stellar radius. For each identified radial pulsation mode (the fundamental and any overtone modes), a frequency mask was applied to the Fourier transform to isolate the spectral region corresponding to the mode of interest. To extract a representative radial profile for the mode, we calculate the mean of the real part of the masked Fourier data. This approach is justified by the prior phase normalisation, which effectively aligns the phase such that the real component predominantly captures the oscillatory behaviour. Averaging along the frequency dimension yields a 1D radial profile, in which the sign indicates the phase state relative to the stellar reference: positive values correspond to regions where the pulsation is in-phase, while negative values indicate antiphase behaviour. Radial nodes, defined as the locations where the pulsation transitions between in-phase and antiphase, are identified by detecting sign changes in this profile. Examples of these phase-normalised amplitude profiles for the first and second overtone pulsation modes of a representative model (st31gm01n003) are shown in Fig. 1. The signal extending beyond the stellar radius indicates outward-travelling waves in the form of shocks (as discussed in Freytag 2017; Ahmad et al. 2023).
Figure 3: Period-luminosity plots for the models in our grid, with each panel representing models of different masses: $1\,\mathrm{M}_{\odot}$, $0.8\,\mathrm{M}_{\odot}$, and $0.7\,\mathrm{M}_{\odot}$ from left to right. The colour bar indicates the summed power of the radial, dipole, and quadrupole velocity power spectra at the stellar surface.
@@ 3.1 Period-luminosity relation
[…]
@@
Figure 5: Ratio of the fundamental to first overtone periods plotted against the fundamental mode period. The colour represents the fraction of power of the fundamental mode, in the power spectrum of the radial velocity, on the stellar surface.
@@ 3.1.1 Comparison with 1D models
The differences between periods derived from our 3D models and those predicted by 1D models, as shown in Fig. 4, are particularly noticeable in the high-luminosity regime for the fundamental mode. These discrepancies can be partially attributed to 3D non-linear effects, but capturing them in 1D models would require more detailed comparisons and parameterisation efforts. In terms of amplitudes, our most Mira-like model has a bolometric amplitude of $\Delta m_{\mathrm{bol}}\approx 0.2\,\mathrm{mag}$, significantly lower than the typical amplitudes of Mira stars observed by Whitelock et al. (2000), which exceed $\Delta m_{\mathrm{bol}}\approx 0.5\,\mathrm{mag}$. In contrast, the amplitudes predicted by Trabucchi et al. (2021) tend to be higher, with the amplitudes being sensitive to the turbulent viscosity parameter. These differences highlight the challenges in reproducing observed amplitudes across different modelling approaches. The interplay between convection and pulsation is evident in the pulsation amplitudes of our AGB models. This interaction is further explored in Sect. 4.3.
@@ 3.2 Radial modes and nodes
We examined the ratio between the fundamental and first overtone modes, in the radial mode, as shown in Fig. 5. We derive that where the ratio lies between $1.7-2.0$, the models in this range corresponds to models with stellar luminosities of approximately between $3.4\leq\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.6$. Above this luminosity range, the period ratios increase significantly. Although overtone modes may still be detectable in this regime, their reliability diminishes, as indicated by their minimal power fraction in the FFT power spectrum, where the fundamental mode dominates. Furthermore, the fundamental mode exhibits a relatively larger spread in its signal. This, combined with weaker and more dispersed signals from overtone modes, results in greater uncertainty in the calculated period ratios.
@@
To highlight the presence of radial nodes in our models and complementary to Fig. 1, Fig. 6 presents the power spectra of three representative models, all corresponding to $1\,\mathrm{M}_{\odot}$ stars exhibiting distinct self-excited pulsation behaviours: a Mira-like star (model st28gm06n073, hereafter model A), a star with irregular pulsations (model st28gm05n053, hereafter model B), and a star dominated by strong overtone pulsations (model st31gm01n003, hereafter model C). These models have average stellar luminosities of $6000$, $5000$, and $2500\,\mathrm{L}_{\odot}$, and effective temperatures of $2800$, $2800$, and $3100\,\mathrm{K}$, respectively. The properties of their simulated luminosities are examined in Sect. 3.5. In the power spectra, nodes appear as minima in power, with the node corresponding to the first overtone in model C being particularly prominent.

### v2

@@ 1 Introduction
Mode switching, where a star changes its dominant pulsation mode (defined as the mode with the highest amplitude), is another phenomenon closely tied to stellar evolution. This transition is expected as stars evolve across different P-L sequences, with observational evidence showing that red giants can switch between fundamental and overtone modes (Kiss et al. 2000; Hinkle et al. 2002). The combined amplitude of decaying higher-overtone modes and emerging lower-overtone modes may be a key factor in triggering more complex pulsation behaviours, such as the long secondary period (Pawlak 2021). The transition from small-amplitude red giants to semiregular variables (SRVs), associated with the onset of self-excited pulsations, is linked to changes in the stellar envelope (Trabucchi & Pastorelli 2025). These changes in the pulsation modes are influenced by factors such as metallicity, hydrogen content, and the treatment of turbulent viscosity in the models. The sensitivity of the transitions of the pulsation modes to the stellar envelope composition provides critical constraints for theoretical models, particularly in understanding the evolutionary pathways of long-period variables (LPVs) (Trabucchi & Pastorelli 2025). Mode switching offers insights into the internal structure of stars and the mechanisms governing the stability of pulsation modes, with implications for stellar feedback and the calibration of P-L relations.
@@ 2.2 Advances over previous generation of models
A significant improvement in this new grid is the adjustment and consistent application of the gravitational smoothing parameter, $r_{0}$ (see Ahmad et al. 2023), across the new model grid. This was set to $7\,\%$ of the stellar radius ($\mathrm{R}_{\star}$), in comparison to between $20-25\,\%$ of $\mathrm{R}_{\star}$ in our previous models. We define $\mathrm{R}_{\star}$, as the point of minimum entropy, based on prior investigations into the optimal definition of this parameter in our models (Ahmad et al. 2023). The $r_{0}$ parameter prescribes the smoothing of the stellar gravitational potential to account for the unresolved compact core, where most of the stellar mass resides, given that we have set the core mass of $0.6\,\mathrm{M}_{\odot}$. Within $2\,r_{0}$, energy is injected at a constant rate according to the desired luminosity. The size of the computational domain and the number of grid points were chosen to keep the number of cells per surface pressure height similar and to account for the different extensions of the atmospheres of models with different surface gravities.
We designate three representative models in this article to discuss specific methods and results. All three have a current stellar mass of $1\,\mathrm{M}_{\odot}$. Their fundamental parameters are listed in Table 1. Chosen to illustrate distinct self-excited pulsation behaviours, the models are referred to by the following letters: a Mira-like star (model A), a star with irregular pulsations (model B), and a star with strong overtone pulsations (model C).
Table 1: Stellar parameters of the representative models used in this article. All three models are of current stellar mass $1\,M_{\odot}$, and the stellar parameters are temporally-averaged values.
Model Designation $R_{\star}\,[R_{\odot}]$ $T_{\text{eff}}\,[\text{K}]$ $L_{\star}\,[L_{\odot}]$
st28gm06n073 A 321 2824 5914
st28gm05n053 B 293 2846 5088
st31gm01n003 C 167 3135 2436
@@ 2.3 Pulsation period extraction process
[…]
@@
Applying the FFT to the spherically averaged velocity yields FFT spectra as a function of radial distance from the stellar centre, corresponding to layers within the stellar interior. To account for spatial variability across these radial layers, an average power spectrum is computed for the region spanning $80-100\,\%$ of $\mathrm{R}_{\star}$. This region consistently exhibits clear signals of pulsations and their effects (see Ahmad et al. 2023), and the resulting averaged power spectrum is thus used as the representative signal for the subsequent analysis. To enhance the clarity of this signal, a combination of sequential smoothing methods is applied. First, a Savitzky-Golay filter is applied to the spectrum using a fourth-degree polynomial and a symmetric window defined by a cut-off frequency of $10^{-3}$ cycles per day. This step retains local features while removing small-scale fluctuations. Second, the output is convolved with a Gaussian kernel with $\sigma=5$ frequency bins (i.e. five times the frequency resolution of the power spectrum) to suppress residual noise. This combined approach has proven effective in retaining the primary pulsation features by preserving peak structures while minimising spurious variability. The improvement is illustrated in Fig. 1, where the final smoothed signal more clearly reveals the local minima surrounding each peak. This facilitates the identification of well-defined frequency bounds for regions containing significant signal power. Overall, this ensured reliable identification of pulsation features in Fourier space, and mitigates non-linear effects such as interference by convection.
Figure 1: The grey curve shows the averaged power spectrum (normalised so that its maximum is unity) of the spherically averaged radial velocity, computed over the radial range $80-100\,\%$ of $\mathrm{R}_{\star}$, for model C. The black curve shows the same signal after applying the two-step smoothing procedure (Savitzky–Golay followed by Gaussian filtering), demonstrating the effectiveness of the method in reducing noise while preserving peak structure. Dotted and dashed vertical lines indicate the derived $f_{\text{mean}}$ values and their corresponding bounds $f_{\textrm{mean}}\pm f_{\textrm{RMS}}$ for the fundamental, first overtone, and second overtone modes, shown in blue, red, and orange, respectively. See main text for further details.
Significant frequencies are identified using an iterative peak-detection procedure inspired by pre-whitening techniques in asteroseismology. At each iteration, the strongest peak in the current power spectrum is located, and an initial frequency range encompassing the peak is defined, by bounding the peak with the nearest local minima on either side of the peak. This region is then refined by computing the weighted mean frequency ($f_{\textrm{mean}}$) and weighted root mean square (RMS) values ($f_{\textrm{RMS}}$) using the power at each frequency bin as weights. We take the frequency region $f_{\textrm{mean}}\pm f_{\textrm{RMS}}$ to be associated with the present pulsation mode. The power contained within this region is normalised by the total spectral power of the original signal, yielding a fractional power that serves as a measure of the mode’s significance. Once characterised, the region is removed (i.e. zeroed) from the original spectrum, and the process is repeated on the residual, including the two-step smoothening process.
Using the procedure described in the preceding paragraph, we identify up to 20 regions. If any of the regions overlap, that is, if their intervals (expanded by the constant factor 1.1 to $f_{\textrm{mean}}\pm 1.1f_{\textrm{RMS}}$ ) intersect, they are merged into a single region. The new bounds are set to the minimum and maximum frequencies across the overlapping intervals. Within this merged region, the $f_{\textrm{mean}}$, $f_{\textrm{RMS}}$ and power fraction are recomputed using the updated window. A justification for this step is seen in Fig. 1: the weaker signal just below $0.006\,\textrm{day}^{-1}$ should be considered as part of the fundamental mode (attributed to non-linear effects, as discussed in Ahmad et al. 2023). But this region would initially be identified as separate from the stronger signal between $0.006-0.008\,\textrm{day}^{-1}$. Merging consolidates these two intervals into a single mode. At the end of this procedure, a minimum of one frequency window remains which correspond to the fundamental mode, and up to four windows may be identified in total, depending on the stellar parameters. These additional windows are associated with overtone modes as seen in Fig. 1 and are discussed further in Sect. 3. For each window, the pulsation period and its spread (could be accounted for as its uncertainty) is then calculated with $P_{\textrm{puls}}=1/f_{\textrm{mean}}$, and its spread $\sigma_{\textrm{puls}}=P_{\textrm{puls}}(f_{\textrm{RMS}}/f_{\textrm{mean}})$.
@@
The method outlined in this section successfully identifies various pulsation modes across different radial orders. The overtone order is determined systematically: the fundamental mode corresponds to the lowest frequency, the first overtone to the second-lowest, and so on. Non-radial modes are distinguished by the velocity components used in the power spectrum analysis. The derived pulsation frequencies and their spreads are used to calculate the fraction of power within each mode’s frequency band. Consequently, we now have periods, frequency spreads, and power fractions for each radial order and degree, allowing for direct comparisons between each of the derived modes. To evaluate luminosity variations, we calculate the spherically averaged luminosity at the outer boundary of the computational domain. While spherical averaging may retain some imprints of non-radial modes, these are less distinct and not reliably distinguishable. However, radial orders remain well separated and can be confidently analysed.
@@ 2.4 Radial node extraction process
Figure 2: The phase-normalised amplitude derived from the Fourier transform of the quantity $\sqrt{\rho}v_{\mathrm{radial}}$, in the frequency bands of the first and second overtones (for model C), in the top and bottom row, respectively.
We further utilise the velocity data within the star to identify radial nodes when an overtone is detected using the process described above. First, we take the spherically averaged quantity $\sqrt{\rho}v_{\mathrm{radial}}(r,t)$ (see Sect. 2.3) as an effective measure of density variations within the star. The complex Fourier transform of this quantity is calculated, now as a function of radial distance and frequency, and subsequently used in the following steps.
The node extraction process begins with phase alignment, where the phase of the complex Fourier-transformed signal is shifted such that the phase at the stellar radius is set to zero. This is implemented by dividing each radial Fourier profile by the complex phase at the radius, leaving only the relative phase differences across the radial layers. For each identified radial mode (fundamental or overtone), a frequency mask defined by $f_{\textrm{mean}}\pm f_{\textrm{RMS}}$ was applied to isolate the corresponding spectral window, which was derived in the step outlined in Sect. 2.3. This ensures that only the signal associated with the mode of interest is considered in the subsequent steps.
At each frequency within the masked range, the Fourier-transformed signal provides a radial profile of the oscillation amplitude. These profiles are then averaged across the frequency window to produce a single, representative radial profile for the mode. We refer to this averaged signal as a phase-normalised amplitude. This averaged profile is still complex-valued, and its real part predominantly captures the amplitude of the standing wave of the pulsation mode. We extract the real part of this profile and interpret the sign to determine the relative phase of the pulsation at each radius: positive values indicate in-phase motion with the stellar surface, while negative values indicate anti-phase behaviour. Radial nodes are then identified as the locations where this profile crosses zero, marking phase transitions. Figure 2 shows examples of these phase-aligned amplitude profiles for the first and second overtone modes of model C. The non-zero signal extending beyond the stellar radius reflects travelling waves that turn into outward-propagating shocks, as discussed by Freytag (2017) and Ahmad et al. (2023).
@@ 3 Results
Figure 4: Period-luminosity plots for the models in our grid, with each panel representing models of different masses: $1\,\mathrm{M}_{\odot}$, $0.8\,\mathrm{M}_{\odot}$, and $0.7\,\mathrm{M}_{\odot}$ from left to right. The colour bar indicates the summed power of the radial, dipole, and quadrupole velocity power spectra at the stellar surface; with darker colours representing higher values. Note that the rows are artificially slanted by a constant factor to improve visual separation and highlight the progression of spectral features.
@@ 3.1 Period-luminosity relation
Overall, the fundamental, first, and second radial overtones can be distinguished clearly, with a smaller scatter within each ridge in Fig. 4. As the pulsation period decreases with increasing radial order, the corresponding ridges in the P-L diagram can be broadly associated with the observed sequences: the fundamental mode ($n=0$) with sequence C, the first overtone ($n=1$) spanning sequences C’ and B, and the second overtone ($n=2$) primarily aligning with sequence A (e.g. Wood et al. 1999; Takayama et al. 2013; Wood 2015; Trabucchi et al. 2019). For the fundamental mode, our previous models have been shown to be in agreement with the P-L relation derived from observations of Miras pulsating in the fundamental mode, after applying bolometric corrections (Ahmad et al. 2023). The fundamental pulsation periods in the current grid, for stars with similar stellar parameters between the previous and current grids, remain consistent despite adjustments to the model settings described in Sect. 2.2. Consequently, we do not recompare the P-L relation with observations of Miras in this study.
@@
Figure 6: Ratio of the fundamental to first overtone periods plotted against the fundamental mode period. The colour represents the fraction of power of the fundamental mode, in the power spectrum of the radial velocity, on the stellar surface. The smaller circles are stars identified as either SRVs or Miras from the OGLE-III catalogue (see text for details).
@@ 3.1.1 Comparison with 1D models
The differences between the fundamental mode periods derived from our 3D models and those predicted by 1D models, as shown in Fig. 5, are most pronounced at both the high- and low-luminosity ends of the model sequence. At high luminosities, the 3D periods tend to be longer than their 1D counterparts, likely reflecting non-linear effects. Conversely, in the low-luminosity regime, the 3D periods are systematically shorter than those predicted by 1D models. The discrepancy likely stems from structural differences between the 1D and 3D models (e.g. the stellar radii that each predicts). Further elucidating these differences would require more detailed comparisons and parameterisation efforts.
In terms of amplitudes, our most Mira-like model has a bolometric amplitude of $\Delta m_{\mathrm{bol}}\approx 0.2\,\mathrm{mag}$, significantly lower than the typical amplitudes of Mira stars observed by Whitelock et al. (2000), which exceed $\Delta m_{\mathrm{bol}}\approx 0.5\,\mathrm{mag}$. In contrast, the amplitudes predicted by Trabucchi et al. (2021) tend to be higher, with the amplitudes being sensitive to the turbulent viscosity parameter. These differences highlight the challenges in reproducing observed amplitudes across different modelling approaches. The interplay between convection and pulsation is evident in the pulsation amplitudes of our AGB models. This interaction is further explored in Sect. 4.3.
@@ 3.2 Radial modes and nodes
We examined the period ratio between the fundamental and first overtone modes, in the radial mode, as shown in Fig. 6. For our models, we derive that where the ratio lies between $1.7-2.0$, the models in this range corresponds to models with stellar luminosities of approximately between $3.4\leq\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.6$. Above this luminosity range, the period ratios increase significantly. Although overtone modes may still be detectable in this regime, their reliability diminishes, as indicated by their minimal power fraction in the FFT power spectrum, where the fundamental mode dominates. Furthermore, the fundamental mode exhibits a relatively larger spread in its signal. This, combined with weaker and more dispersed signals from overtone modes, results in greater uncertainty in the calculated period ratios.
@@
Additionally, in Fig.6, we include SRVs and Miras from the OGLE-III catalogue of LPVs in the Large Magellanic Cloud (LMC) (Soszyński et al. 2009). The aim is to visualise the observed spread in the ratio of the fundamental to first overtone period, and to compare this with our models. To this end, we derived a subset of 3429 sources (3201 SRVs and 228 Miras) by restricting the OGLE-III dataset to those that fall within the axis limits of Fig. 6. While the OGLE-III catalogue provides amplitude-based mode assignments (as analysed in Wood 2015), we instead order the periods for each star by length, assuming the longest period corresponds to the fundamental mode, followed by the first overtone. This comparison highlights that the spread in the period ratios produced by our models aligns well with the observed distribution of LPVs, lending further confidence to the physical realism of the simulations. We note, however, that the assumption of mode ordering by period introduces some uncertainty: in particular, the observed spread may partially reflect differences in angular degree, such as the fundamental mode being radial while the overtone corresponds to a non-radial mode (e.g. dipole). Nonetheless, the similarity in the spread of the period ratios reinforces the consistency between the models and observations.
To highlight the presence of radial nodes in our models and complementary to Fig. 2, Fig. 7 presents the power spectra of three representative models, all corresponding to $1\,\mathrm{M}_{\odot}$ stars exhibiting distinct self-excited pulsation behaviours as described in Sect. 2.2. The properties of their simulated luminosities are examined in Sect. 3.5. In the power spectra presented in Fig. 7, nodes appear as minima in power, with the node corresponding to the first overtone in model C being particularly prominent.